OpenStudy (anonymous):

solve each equation. if the equation is an identity or if it has no solution, write identity or no solution. 2(g-2) - 4=2(g-3)

OpenStudy (anonymous):

what do you think ? \[2(g-2)-4=2(g-2-2)=2(g-4)\]